A child is pushing a merry-go-round. The angle through which the merry-go-round has turned varies with time according to ?(t)=?t+?t3, where ?= 0.383 rad/s and ?= 1.40
Answer #1

PART A:
angular velocity = omega(t)
in order to find angular velocity, differentiate theta(t) given in the question.
answ:
omega(t) = (gamma)+(3*beta*t^2) rad/s

PART B:
omega_0 occur when t=0
from equation in PART B: omega(t) = (gamma)+3(beta)*t^2
substitute gamma = 0.383, beta = 0.0140 and t=0
answ:
omega_0 : 0.383 + (3*0.0140*0^2) = 0.383 rad/s

PART C:
the working is as in PART B
omega(5) = (gamma)+3(beta)*t^2 = 0.383 + (3*0.0140*4.95^2) = 1.412 rad/s

PART D:
the formula of average angular velocity for the time interval t = 0 to t = 4.95 seconds is : (theta(4.95) - theta(0)) / ( 4.95-0 )
answ:
(0.383(4.95) + (0.014*(4.95^3)) - 0) / (4.95-0) = 0.726 rad/s
Answer #2

PART A:
angular velocity = omega(t)
in order to find angular velocity, differentiate theta(t) given in the question.
answ:
omega(t) = (gamma)+(3*beta*t^2) rad/s

PART B:
omega_0 occur when t=0
from equation in PART B: omega(t) = (gamma)+3(beta)*t^2
substitute gamma = 0.383, beta = 0.0140 and t=0
answ:
omega_0 : 0.383 + (3*0.0140*0^2) = 0.383 rad/s

PART C:
the working is as in PART B
omega(4.95) = (gamma)+3(beta)*t^2 = 0.383 + (3*0.0140*4.95^2) = 1.412 rad/s

PART D:
the formula of average angular velocity for the time interval t = 0 to t = 4.95 seconds is : (theta(4.95) - theta(0)) / ( 4.95-0 )